# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 6.1. Limitations
L5 Code and Vulnerability Representation
In most surveyed studies, the code to be analyzed is provided as raw code or embedded directly within the prompt during inference (cf. Figure 2). However, models frequently rely on superficial textual patterns rather than capturing the underlying semantic structures of vulnerabilities (Ding et al., 2025; Zhao et al., 2023; Li et al., 2025a). As a result, model performance degrades under semantic-preserving code transformations, such as variable renaming, code reordering, or formatting changes, highlighting a lack of robustness and true semantic understanding (Ni et al., 2024a; Risse and B”̈ohme, 2024; Qi et al., 2024; Hort et al., 2025; Panichella, 2025; Wang et al., 2025b). While contrastive learning is designed to produce more discriminative embeddings, models still often fail to draw the classification boundary correctly and identify the vulnerable patterns (Ding et al., 2025).
L6 Model Interpretability and Explainability
Current LLM-based vulnerability detection approaches offer limited insight into why a model classifies the input code as vulnerable or non-vulnerable. Even in studies that use CoT prompting or explicitly investigate reasoning as an additional objective, issues with hallucinated or misleading justifications are commonly reported (Espinha Gasiba et al., 2024; Noever, 2023). Similarly, studies implementing self-verification report instances of inaccurate corrections, self-contradictions, and hallucinated justifications (Yu et al., 2024). These issues undermine the trustworthiness of model outputs and hinder their integration in development workflows where interpretability and reliability are critical.
L7 Integration into Pipelines and Workflows
Despite growing interest, the integration of LLM-based vulnerability detection into practical development workflows remains underexplored. Only a few studies investigate real-world applicability through prototype tools or user studies (Fu et al., 2024; Steenhoek et al., 2025b), though such efforts are essential for assessing feasibility and trust in practical settings. Generalization to realistic codebases also remains challenging, with models often underperforming outside curated benchmarks (Chakraborty et al., 2024).
@@ 6.2. Actionable Insights and Future Research Opportunities
L1 Dataset and Detection Granularity
To move beyond the limitations of function-level detection, future work should focus on context-enhanced vulnerability detection, including project- and repository-level analysis (Zhou et al., 2024a; Wen et al., 2024b; Li et al., 2025d; Guo et al., 2025). Initial steps in this direction include augmenting function-level samples with additional context such as surrounding code lines (Halder et al., 2025), function arguments, external functions, type definitions, global variables, environmental constraints (Risse et al., 2025), as well as dependency and execution flows (Zheng et al., 2025). For example, Ahmed et al. (Ahmed et al., 2025) use GPT-4 to identify the required contexts for a given vulnerability in a function, but still find limitations in pinpointing vulnerable statements and their root causes in complex real-world code. Li et al. (Li et al., 2025e) demonstrate that incorporating execution and data context improves model performance and reasoning quality, emphasizing that the challenge lies in precise, context-aware vulnerability reasoning. Similarly, Yang et al. (Yang et al., 2025b) propose a program analysis-based approach that abstracts complex function calls into primitive API representations to enrich contextual understanding. Further investigating such context-enhancements is essential for capturing dependencies that span multiple functions or files, enabling the detection of more complex vulnerability types that are otherwise missed in isolated, function-level analysis.
@@
L2 Dataset Labeling and Quality
To address current limitations in dataset construction, future work should focus on building real-world vulnerability datasets with improved label quality and balanced coverage across CWE types. Label quality can be improved through two main strategies: (1) refining existing datasets via targeted manual labeling to correct and verify automatically generated labels (He and Vechev, 2023), and (2) advancing automated labeling methods, such as VFCs, to reduce noise and improve labeling reliability (Ding et al., 2024a; Li et al., 2025f; Ahmed et al., 2025; Gao et al., 2025; Lu et al., 2025). Balanced coverage across frequent and less common vulnerability types can be achieved, e.g., by using LLMs to generate vulnerable samples for specific vulnerability types (Yong et al., 2025; Bappy et al., 2025; Shahzad et al., 2025).
@@
L3 Study Evaluation and Comparability
@@
To enhance reproducibility and comparability across studies, future research should prioritize standardized evaluation protocols. Specifically, we recommend the use of datasets with documented train/validation/test splits and careful alignment with datasets used in related studies to ensure meaningful comparisons within the same use case. We provide the mappings of datasets to surveyed use cases in the artifacts (Kaniewski et al., 2025).
@@
To assess generalization and maximize comparability with prior works, models should be evaluated on multiple datasets, combining recent datasets such as PrimeVul (Ding et al., 2024a) or MegaVul (Ni et al., 2024b) with established benchmark datasets such as Devign (Zhou et al., 2019) or Big-Vul (Fan et al., 2020). All pre-processing and filtering steps must be documented to enable replication. In support of open science (Nong et al., 2023), sharing of code, evaluation scripts, and model checkpoints is encouraged.
L4 Up-to-Date Vulnerability Knowledge
@@
Given the broad and continuously growing number of CVEs, ensuring that models remain informed about newly discovered vulnerabilities is crucial for maintaining relevance in practical scenarios. To address the limitation of static datasets quickly becoming outdated, future work should focus on continuously maintained and automatically extensible datasets, such as CVEfixes (Bhandari et al., 2021), which employs an automated pipeline for updates. Advanced adaptation techniques also offer promising pathways to bridge this knowledge gap. RAG enables LLMs to access external, up-to-date knowledge at inference time. Continual learning can further support adaptability by integrating new vulnerability knowledge while preserving performance on previously seen vulnerability types. However, both techniques have seen limited adoption in the context of software vulnerability detection (cf. Figure 2). Further research should explore their effectiveness, especially in handling large and heterogeneous vulnerability information. For RAG, this includes investigating how different knowledge base representations, e.g., structured CWE hierarchy graphs or abstract function descriptions, affect retrieval quality and downstream detection accuracy.
L5 Code and Vulnerability Representation
@@
To address the need for robust representations that capture vulnerability semantics, future work should further focus on structure-aware input representations. Directions include graph-based representations, such as control and data flow graphs, (LLM-generated) abstract descriptions that capture semantics and code behavior (Wu et al., 2025; Du et al., 2025), and multi-modal inputs, e.g., combining API abstractions, data flow graphs, and natural language documentation (Yang et al., 2025b). Understanding how different representations complement each other can enhance vulnerability detection capabilities. Ideally, such representations should be language-agnostic to support generalization across diverse programming languages (cf. multi-language studies (Dozono et al., 2024; Zhang et al., 2025b)). In addition, hybrid architectures should be further explored. GNNs, in particular, have demonstrated potential for capturing code semantics and structural dependencies (Chang et al., 2024; Islam et al., 2023, 2024a; Jiang et al., 2024b; JianJie and Le, 2023; Kong et al., 2024; Liu et al., 2025c; Ni et al., 2023a; Yang et al., 2024a; Zheng et al., 2024). Future work should also evaluate the robustness of different representation strategies against syntactic perturbations and domain shifts, such as LLM-generated or refactored code, to better assess real-world generalization capabilities.
L6 Model Interpretability and Explainability
@@
To build trust in LLM-based vulnerability detection systems, future research should go beyond standard evaluation metrics (e.g., accuracy, recall, precision, F1) and incorporate dedicated metrics that assess the trustworthiness and consistency of model-generated explanations. As LLMs increasingly take on roles in security workflows, their predictions must not only be correct but also justifiable and transparent to developers. Justification includes distinguishing between correct predictions made for the wrong reasons (spurious correlations) and those grounded in valid vulnerability semantics (Risse et al., 2025).
@@
Interpretability may further be enhanced through structured explanation formats, such as vulnerability propagation paths or flow annotations. These structured rationales offer greater reliability than natural language justifications and can be cross-checked against program logic or expert knowledge. Further, integrating RAG techniques with structured vulnerability knowledge, such as CWE hierarchies or exploit chains, could improve reasoning by anchoring the model’s output in explicit and verifiable context. This approach may not only strengthen the factual grounding of explanations but also aligns with the information needs of developers and auditors in practice.
L7 Integration into Pipelines and Workflows
@@
Agentic systems and reasoning-driven workflows offer promising pathways to autonomously assist developers through tasks such as self-assessment, explanation generation, or iterative refinement. For example, Farr et al. (Farr et al., 2025) propose an expert-in-the-loop system to route code into the three categories ”automatic quarantine”, ”deployment clearance”, or ”manual review”, based on model confidence scores, helping prioritize expert attention where most needed.
To further improve generalization to practical vulnerability remediation workflows, approaches should move towards CWE-type classification. Promising directions include the development of type-specific classifiers that align with the CWE hierarchy and research views, cf. (Atiiq et al., 2024b; Cao et al., 2024; Deng et al., 2024), enabling more actionable insights and effective prioritization.
@@ 7. Conclusion
In this systematic literature review, we analyzed 227 studies on LLM-based software vulnerability detection published between January 2020 and June 2025. To structure this rapidly evolving field, we introduced a comprehensive taxonomy that covers task formulation, i.e., the classification task presented to the LLM; input representation, i.e., how code and context are provided to the LLM; system architecture, differentiating between LLM-centric and hybrid approaches; and model adaptation techniques, including prompt engineering, fine-tuning, and learning paradigms. The analysis shows that most studies use simple approaches, such as binary code classification (vulnerable/non-vulnerable) and general-purpose LLMs. Common techniques include full-parameter fine-tuning and zero-shot prompting. More advanced techniques such as parameter-efficient fine-tuning, continual learning, retrieval-augmented generation, and agentic workflows are only beginning to emerge. We further analyzed the datasets used for training and evaluation, extending the taxonomy to include dataset type, granularity, source, and labeling. We investigated commonly used datasets with respect to their realism, vulnerability coverage, diversity, and usage trends. Despite notable progress in the field, we identified several key limitations that hinder practical adoption of LLMs in software vulnerability detection, including limited detection granularity, limited robustness in code representations, and outdated vulnerability knowledge. In particular, dataset-related limitations, such as low vulnerability type diversity and class imbalance, pose challenges to generalization and cross-study comparability. To address these limitations, we outlined actionable research directions, such as advancing structure-aware and language-agnostic input representations, aligning datasets across use cases to enable cross-study comparison, and adopting RAG and continual learning techniques to improve adaptability. By mapping existing studies, identifying open challenges, and proposing future research directions, this review aims to guide researchers and practitioners in selecting techniques and datasets as well as support more comparable and reproducible research; ultimately advancing the development of reliable, generalizable, and practically applicable LLM-based vulnerability detection systems.
@@ Acknowledgements.
- Akli et al. (2025) Amal Akli, Maxime Cordy, Mike Papadakis, and Yves Le Traon. 2025. AutoAdapt: On the Application of AutoML for Parameter-Efficient Fine-Tuning of Pre-Trained Code Models. ACM Transactions on Software Engineering and Methodology (TOSEM) Just Accepted, Just Accepted (2025). doi:10.1145/3734867
- Allamanis et al. (2018) Miltiadis Allamanis, Earl T. Barr, Premkumar Devanbu, and Charles Sutton. 2018. A Survey of Machine Learning for Big Code and Naturalness. ACM Computing Surveys (CSUR) 51, 4, Article 81 (2018), 37 pages. doi:10.1145/3212695
@@
- Cao et al. (2024) Di Cao, Yong Liao, and Xiuwei Shang. 2024. RealVul: Can We Detect Vulnerabilities in Web Applications with LLM? arXiv:2410.07573
@@
- Chen et al. (2025) Siyu Chen, Jiongyi Yang, Xiang Chen, Menglin Zheng, Minnan Wei, and Xiaolin Ju. 2025. Improving Vulnerability Type Prediction and Line-Level Detection via Adversarial Training-based Data Augmentation and Multi-Task Learning. arXiv:2506.23534
@@
- Dai et al. (2025) Shih-Chieh Dai, Jun Xu, and Guanhong Tao. 2025. A Comprehensive Study of LLM Secure Code Generation. arXiv:2503.15554
@@
- Du et al. (2024) Xiaohu Du, Ming Wen, Jiahao Zhu, Zifan Xie, Bin Ji, Huijun Liu, Xuanhua Shi, and Hai Jin. 2024. Generalization-Enhanced Code Vulnerability Detection via Multi-Task Instruction Fine-Tuning. arXiv:2406.03718
- Du et al. (2025) Xueying Du, Geng Zheng, Kaixin Wang, Yi Zou, Yujia Wang, Wentai Deng, Jiayi Feng, Mingwei Liu, Bihuan Chen, Xin Peng, Tao Ma, and Yiling Lou. 2025. Vul-RAG: Enhancing LLM-based Vulnerability Detection via Knowledge-level RAG. arXiv:2406.11147
@@
- Fu et al. (2025) Yujia Fu, Peng Liang, Amjed Tahir, Zengyang Li, Mojtaba Shahin, Jiaxin Yu, and Jinfu Chen. 2025. Security Weaknesses of Copilot-Generated Code in GitHub Projects: An Empirical Study. arXiv:2310.02059
- Gao et al. (2023b) Shuzheng Gao, Hongyu Zhang, Cuiyun Gao, and Chaozheng Wang. 2023b. Keeping Pace with Ever-Increasing Data: Towards Continual Learning of Code Intelligence Models. In International Conference on Software Engineering (ICSE). IEEE/ACM, Melbourne, Australia, 30–42. doi:10.1109/ICSE48619.2023.00015
- Gao et al. (2023a) Zeyu Gao, Hao Wang, Yuchen Zhou, Wenyu Zhu, and Chao Zhang. 2023a. How Far Have We Gone in Vulnerability Detection Using Large Language Models. arXiv:2311.12420
@@
- Ghafoor Hussain et al. (2025) Rida Ghafoor Hussain, Kin-Choong Yow, and Marco Gori. 2025. Leveraging an Enhanced CodeBERT-Based Model for Multiclass Software Defect Prediction via Defect Classification. IEEE Access 13 (2025), 24383–24397. doi:10.1109/ACCESS.2024.3525069

### v2

@@ 6.1.4. L4 Up-to-Date Vulnerability Knowledge
6.1.5. L5 Code and Vulnerability Representation
In most surveyed studies, the code to be analyzed is provided as raw code or embedded directly within the prompt during inference. However, models frequently rely on superficial textual patterns rather than capturing the underlying semantic structures of vulnerabilities (Ding et al., 2025; Zhao et al., 2023; Li et al., 2025a). As a result, model performance degrades under semantic-preserving code transformations, such as variable renaming, code reordering, or formatting changes, highlighting a lack of robustness and true semantic understanding (Ni et al., 2025; Risse and B”̈ohme, 2024; Qi et al., 2024; Hort et al., 2025; Panichella, 2025; Wang et al., 2025b). While contrastive learning is designed to produce more discriminative embeddings, models still often fail to draw the classification boundary correctly and identify the vulnerable patterns (Ding et al., 2025).
6.1.6. L6 Model Interpretability and Explainability
Current LLM-based vulnerability detection approaches offer limited insight into why a model classifies the input code as vulnerable or non-vulnerable. Even in studies that use CoT prompting or explicitly investigate reasoning, issues with hallucinated or misleading justifications are commonly reported (Espinha Gasiba et al., 2024; Noever, 2023). Similarly, studies implementing verification report instances of inaccurate corrections, self-contradictions, and hallucinated justifications (Yu et al., 2024). These issues undermine the trustworthiness of model outputs and hinder their integration in practical development workflows where interpretability and reliability are critical.
6.1.7. L7 Integration into Pipelines and Workflows
Despite growing interest, the integration of LLM-based vulnerability detection into practical development workflows remains underexplored. Only a few studies investigate real-world applicability through prototype tools or user studies (e.g., (Fu et al., 2024; Steenhoek et al., 2025b)), though such efforts are essential for assessing feasibility and trust in practical settings. Generalization to realistic codebases also remains challenging, with models often underperforming outside curated benchmarks (Chakraborty et al., 2024).
@@ 6.2. Actionable Insights and Future Research Opportunities
6.2.1. L1 Dataset and Detection Granularity
To move beyond the limitations of function-level detection, future work should focus on context-enhanced vulnerability detection, including project- and repository-level analysis (Zhou et al., 2024a; Wen et al., 2024b; Li et al., 2025f; Guo et al., 2025). Initial steps in this direction include augmenting function-level samples with additional context, such as surrounding code lines (Halder et al., 2025), function arguments, external functions, type definitions, global variables, environmental constraints (Risse et al., 2025), as well as dependency and execution flows (Zheng et al., 2025). For example, Ahmed et al. (Ahmed et al., 2025) use GPT-4 to identify the required contexts for a given vulnerability in a function, but still find limitations in pinpointing vulnerable statements and their root causes in complex real-world code. Li et al. (Li et al., 2025i) demonstrate that incorporating execution and data context improves model performance and reasoning quality, emphasizing that the challenge lies in precise, context-aware vulnerability reasoning. Similarly, Yang et al. (Yang et al., 2025b) propose a program analysis-based approach that abstracts complex function calls into primitive API representations to enrich contextual understanding. Further investigating such context-enhancements is essential for capturing dependencies that span multiple functions or files, enabling the detection of more complex vulnerability types that are otherwise missed in isolated, function-level analysis.
@@ 6.2.1. L1 Dataset and Detection Granularity
6.2.2. L2 Dataset Labeling and Quality
To address current limitations in dataset construction, future work should focus on building real-world vulnerability datasets with improved label quality and balanced coverage across CWE types. Label quality can be improved through two main strategies: (1) refining existing datasets via targeted manual labeling to correct and verify automatically generated labels (He and Vechev, 2023), and (2) advancing automated labeling methods, such as VFCs, to reduce noise and improve labeling reliability (Ding et al., 2024a; Li et al., 2025k; Ahmed et al., 2025; Gao et al., 2025; Lu et al., 2025). Balanced coverage across frequent and less common vulnerability types can be achieved, e.g., by using LLMs to generate vulnerable samples for specific vulnerability types (Yong et al., 2025; Bappy et al., 2025; Shahzad et al., 2025).
Complementary to improved context, there is a critical need for further reasoning-oriented benchmarks (Sun et al., 2025b) to close the gap in ground truth. Future dataset creation should prioritize pairing vulnerable code with verified, step-by-step explanations of the root cause and exploitation logic. Establishing such reasoning ground truths is a prerequisite for advancing from simple pattern matching to genuine semantic understanding.
@@ 6.2.2. L2 Dataset Labeling and Quality
6.2.3. L3 Study Evaluation and Comparability
@@
To enhance reproducibility and comparability across studies, future research should prioritize standardized evaluation protocols. Specifically, we recommend the use of datasets with documented train/validation/test splits (or even closed evaluation of test splits) and careful alignment with datasets used in related studies to ensure meaningful comparisons within the same use case. We provide the mappings of datasets to surveyed use cases in the artifacts (Kaniewski et al., 2025).
@@
To assess generalization and maximize comparability with prior works, models should be evaluated on multiple datasets, combining recent datasets such as PrimeVul (Ding et al., 2024a) or MegaVul (Ni et al., 2024) with established benchmark datasets such as Devign (Zhou et al., 2019) or Big-Vul (Fan et al., 2020). All pre-processing and filtering steps must be documented to enable replication. In support of open science (Nong et al., 2023), sharing of code, evaluation scripts, and model checkpoints is encouraged.
6.2.4. L4 Up-to-Date Vulnerability Knowledge
@@
Given the broad and continuously growing number of CVEs, ensuring that models remain informed about newly discovered vulnerabilities is crucial for maintaining relevance in practical scenarios. To address the limitation of static datasets quickly becoming outdated, future work should focus on continuously maintained and automatically extensible datasets, such as CVEfixes (Bhandari et al., 2021), which employs an automated pipeline for updates. Advanced adaptation techniques also offer promising pathways to bridge this knowledge gap. RAG enables LLMs to access external, up-to-date knowledge at inference time. Continual learning can further support adaptability by integrating new vulnerability knowledge while preserving performance on previously seen vulnerability types. However, both techniques have seen limited adoption in the context of software vulnerability detection. Further research should explore their effectiveness, especially in handling large and heterogeneous vulnerability information. For RAG, this includes investigating how different representations, e.g., structured CWE hierarchy graphs or abstract function descriptions, affect retrieval quality and downstream detection accuracy.
6.2.5. L5 Code and Vulnerability Representation
@@
To address the need for robust representations that capture vulnerability semantics, future work should further focus on structure-aware input representations. Directions include graph-based representations, such as control and data flow graphs, (LLM-generated) abstract descriptions that capture semantics and code behavior (Zhang et al., 2025a; Wu et al., 2025; Du et al., 2025a), and multi-modal inputs, e.g., combining API abstractions, data flow graphs, and natural language documentation (Yang et al., 2025b). Understanding how different representations complement each other can enhance vulnerability detection capabilities. Ideally, such representations should be language-agnostic to support generalization across diverse programming languages (cf. multi-language studies (Dozono et al., 2024; Zhang et al., 2025b)). In addition, hybrid architectures should be further explored. GNNs, in particular, have demonstrated potential for capturing code semantics and structural dependencies (Chang et al., 2024; Islam et al., 2023, 2024a; Jiang et al., 2024b; JianJie and Le, 2023; Kong et al., 2024; Liu et al., 2025b; Ni et al., 2023a; Yang et al., 2024a; Zheng et al., 2024). Future work should also evaluate the robustness of different representation strategies against syntactic perturbations and domain shifts, such as LLM-generated or refactored code, to better assess real-world generalization capabilities.
6.2.6. L6 Model Interpretability and Explainability
@@
To build trust in LLM-based vulnerability detection systems, future research should go beyond standard evaluation metrics (e.g., accuracy, recall, precision, F1) and incorporate dedicated metrics that assess the trustworthiness and consistency of model-generated explanations. As LLMs increasingly take on roles in security workflows, their predictions must not only be correct but also justifiable and transparent to developers. Justification includes distinguishing between correct predictions made for the wrong reasons (spurious correlations) and those grounded in valid vulnerability semantics (Risse et al., 2025). In this matter, studies should increasingly look into Explainable AI (XAI) techniques (Takieldeen et al., 2025; Nguyen et al., 2025a; Alam et al., 2025).
@@
Interpretability may further be enhanced through structured explanation formats, such as vulnerability propagation paths or flow annotations. These structured rationales offer greater reliability than natural language justifications and can be cross-checked against program logic or expert knowledge. Further, integrating RAG techniques with structured vulnerability knowledge, such as CWE hierarchies or exploit chains, could improve reasoning by anchoring the model’s output in verifiable context. This approach may not only strengthen the grounding of explanations but also align with the information needs of developers and auditors in practice.
6.2.7. L7 Integration into Pipelines and Workflows
@@
Agentic systems and reasoning-driven workflows offer promising pathways to autonomously assist developers (e.g., expert-in-the-loop system (Farr et al., 2025)) through tasks such as self-assessment, explanation generation, or iterative refinement.
To further improve generalization to practical vulnerability remediation workflows, LLM-based approaches should move towards CWE-type classification and report generation. Promising directions include the development of type-specific classifiers that align with the CWE hierarchy and research views, cf. (Atiiq et al., 2024b; Cao et al., 2024; Deng et al., 2024), enabling more actionable insights and effective prioritization.
A current pathway for industrial adoption lies in synergizing LLMs with traditional program analysis (Li et al., 2024). While LLMs excel at semantic reasoning, they lack the formal correctness guarantees of, e.g., formal verification tools. Pipelines where SAST tools detect candidates and LLMs filter false positives can significantly reduce alert fatigue by applying semantic understanding to noisy static reports. Conversely, workflows where LLMs propose vulnerabilities and static analyzers verify feasibility offer a mechanism to ground probabilistic outputs.
@@
7. Threats to Validity
To ensure transparency of our findings, we acknowledge potential threats to the validity of this SLR along with the mitigation strategies employed.
Paper retrieval omissions. A primary threat to any SLR is the omission of relevant studies due to studies with incomplete titles, abstracts, or varied keyword terminologies. To address this threat, we constructed a broad search string based on manually selected keywords from primary studies as well as keywords from prior SLRs. We combined manual search, automated search, and snowballing to minimize overlooking relevant papers. Despite these efforts, due to the living nature of the field, a static snapshot may miss very recent developments. To address this threat, we employ a continuous search strategy using alerts and maintain a repository to update the study corpus.
Study selection bias. The selection of studies involves subjective judgment, which introduces potential bias. To minimize this threat, we conducted a revision of all crawled papers, strictly applying predefined inclusion and exclusion criteria (cf. Table 2). A specific threat regarding quality arises from our decision not to filter exclusively for high-ranked conferences or journals. We included a wide range of studies to provide a holistic picture of the landscape and identify gaps that might be missing in a narrower scope. To address the quality implications of this decision, we provide an analysis of the distribution of studies across conference and journal ranks as a quality indicator. Additionally, we include preprints from arXiv. While necessary to reflect the latest advancements, preprints have not undergone rigorous peer review. To mitigate this threat, we conducted a manual quality assessment, excluding studies that lacked clear evaluation setups or coherent contributions.
Data extraction and classification. The manual extraction of information and the construction of the taxonomy are subject to researcher bias and interpretation errors. For instance, determining whether a study performs ”reasoning” versus ”description” can be subjective. To mitigate this risk, we aligned our taxonomy with existing definitions where possible (e.g., dataset granularity, PEFT techniques) and refined our categories iteratively during the full-text review. We further support the reliability of our extraction by publicly releasing the full list of categorized studies and their attributes in our repository (Kaniewski et al., 2025).
@@ 8. Conclusion
In this SLR, we analyzed 263 studies on LLM-based software vulnerability detection. To structure this rapidly evolving field, we introduced a comprehensive taxonomy that covers detection task formulation, i.e., the classification and generation tasks presented to the LLM; input representation, i.e., how code and auxiliary information are provided to the LLM; system architecture, differentiating between LLM-centric and hybrid approaches; as well as model adaptation (prompt engineering, fine-tuning, and learning paradigms) and orchestration techniques, e.g., agentic and ensemble systems. The analysis shows that most studies perform binary code classification (vulnerable/non-vulnerable). Common adaptation techniques include full-parameter fine-tuning and zero-shot prompting. More advanced adaptation techniques, learning paradigms, and orchestration designs, such as parameter-efficient fine-tuning, retrieval-augmented generation, continual learning, and agentic workflows, are only beginning to emerge. We further analyzed common datasets used, investigating their realistic nature, diversity in vulnerability types with respect to CWE-1000 pillars, and usage trends. Despite notable progress in the field, we identified several key limitations that hinder practical adoption of LLMs in software vulnerability detection, including limited robustness in code representations, vulnerability data leakage, and limited interpretability and explainability. To address the identified limitations, we outlined actionable research directions, such as advancing structure-aware and language-agnostic input representations, aligning datasets across use cases to enable cross-study comparison and standardized evaluation protocols, adopting retrieval and continual learning techniques to improve adaptability, and fitting approaches into practical development workflow requirements. By mapping existing studies, identifying open challenges, and proposing future research directions, this review aims to guide researchers and practitioners in advancing the development of reliable, generalizable, and practically applicable LLM-based vulnerability detection systems.
@@ Acknowledgements.
- Akli et al. (2025) Amal Akli, Maxime Cordy, Mike Papadakis, and Yves Le Traon. 2025. AutoAdapt: On the Application of AutoML for Parameter-Efficient Fine-Tuning of Pre-Trained Code Models. ACM Transactions on Software Engineering and Methodology (TOSEM) (2025). doi:10.1145/3734867 Just Accepted.
- Alam et al. (2025) Syeda Sadia Alam, Mst Shapna Akter, and Alfredo Cuzzocrea. 2025. Improving Software Security Through a LLM-Based Vulnerability Detection Model. In International Conference on Database and Expert Systems Applications (DEXA). Springer, Bangkok, Thailand, 122–129. doi:10.1007/978-3-032-02049-9_9
- Allamanis et al. (2018) Miltiadis Allamanis, Earl T. Barr, Premkumar Devanbu, and Charles Sutton. 2018. A Survey of Machine Learning for Big Code and Naturalness. Comput. Surveys 51, 4, Article 81 (2018), 37 pages. doi:10.1145/3212695
- Alqarni and Azim (2025) Mansour Alqarni and Akramul Azim. 2025. An Advanced Detection Framework for Embedded System Vulnerabilities. IEEE Access 13 (2025), 159207–159216. doi:10.1109/ACCESS.2025.3607595
- Alrashedy et al. (2024) Kamel Alrashedy, Vincent J. Hellendoorn, and Alessandro Orso. 2024. Learning Defect Prediction from Unrealistic Data. In International Conference on Software Analysis, Evolution and Reengineering (SANER). IEEE, Rovaniemi, Finland, 556–567.
@@
- Aslan et al. (2025) Mert Kaan Aslan, Yunus Emre Alkan, Muhammed Burak Alican, and Özgür Özdemir. 2025. Utilizing Large Programming Language Models on Software Vulnerability Detection. In Innovations in Intelligent Systems and Applications Conference (ASYU). IEEE, Bursa, Turkey, 1–6. doi:10.1109/ASYU67174.2025.11208282
@@
- Cao et al. (2024) Di Cao, Yong Liao, and Xiuwei Shang. 2024. RealVul: Can We Detect Vulnerabilities in Web Applications with LLM?. In Conference on Empirical Methods in Natural Language Processing (EMNLP). ACL, Miami, FL, USA, 8268–8282. doi:10.18653/v1/2024.emnlp-main.472
@@
- Chen et al. (2025b) Siyu Chen, Jiongyi Yang, Xiang Chen, Menglin Zheng, Minnan Wei, and Xiaolin Ju. 2025b. Improving Vulnerability Type Prediction and Line-Level Detection via Adversarial Training-based Data Augmentation and Multi-Task Learning. arXiv:2506.23534
@@
- Chen et al. (2025a) Yinan Chen, Yuan Huang, Xiangping Chen, Pengfei Shen, and Lei Yun. 2025a. GPTVD: vulnerability detection and analysis method based on LLM’s chain of thoughts. Automated Software Engineering (ASE) 33, 1 (2025), 1–23. doi:10.1007/s10515-025-00550-4
- Chen and Zhang (2025) Zhuo Chen and Xuejun Zhang. 2025. VulPr: A Prompt Learning-based Method for Vulnerability Detection. In International Conference on Electronic Information Technology (EIT). IEEE, Valparaiso, IN, USA, 594–597. doi:10.1109/EIT67313.2025.11231886
- Cheng et al. (2025) Zeru Cheng, Yanjing Yang, He Zhang, Lanxin Yang, Jinghao Hu, Jinwei Xu, Bohan Liu, and Haifeng Shen. 2025. Leveraging Self-Paced Learning for Software Vulnerability Detection. arXiv:2511.09212
@@
- Du et al. (2024) Xiaohu Du, Ming Wen, Jiahao Zhu, Zifan Xie, Bin Ji, Huijun Liu, Xuanhua Shi, and Hai Jin. 2024. Generalization-Enhanced Code Vulnerability Detection via Multi-Task Instruction Fine-Tuning. In Findings of the Association for Computational Linguistics (ACL). ACL, Bangkok, Thailand, 10507–10521. doi:10.18653/v1/2024.findings-acl.625
- Du et al. (2025a) Xueying Du, Geng Zheng, Kaixin Wang, Yi Zou, Yujia Wang, Wentai Deng, Jiayi Feng, Mingwei Liu, Bihuan Chen, Xin Peng, Tao Ma, and Yiling Lou. 2025a. Vul-RAG: Enhancing LLM-based Vulnerability Detection via Knowledge-level RAG. arXiv:2406.11147
- Du et al. (2025b) Xiaozhi Du, Yanrong Zhou, and Hongyuan Du. 2025b. DMVL4AVD: A Deep Multi-View Learning Model for Automated Vulnerability Detection. Neural Computing and Applications (Neural Comput. Appl.) 37 (2025), 5873–5889. doi:10.1007/s00521-024-10892-x
@@
- Flicke et al. (2025) Markus Flicke, Glenn Angrabeit, Madhav Iyengar, Vitalii Protsenko, Illia Shakun, Jovan Cicvaric, Bora Kargi, Haoyu He, Lukas Schuler, Lewin Scholz, Kavyanjali Agnihotri, Yong Cao, and Andreas Geiger. 2025. Scholar Inbox: Personalized Paper Recommendations for Scientists. arXiv:2504.08385
@@
- Fu et al. (2025b) Michael Fu, Jirat Pasuksmit, and Chakkrit Tantithamthavorn. 2025b. AI for DevSecOps: A Landscape and Future Opportunities. ACM Transactions on Software Engineering and Methodology (TOSEM) 34, 4, Article 115 (2025), 61 pages. doi:10.1145/3712190
@@
- Fu et al. (2025a) Yujia Fu, Peng Liang, Amjed Tahir, Zengyang Li, Mojtaba Shahin, Jiaxin Yu, and Jinfu Chen. 2025a. Security Weaknesses of Copilot-Generated Code in GitHub Projects: An Empirical Study. doi:10.1145/3716848
- Gajjar et al. (2025a) Jugal Gajjar, Kaustik Ranaware, and Kamalasankari Subramaniakuppusamy. 2025a. Bridging Semantics & Structure for Software Vulnerability Detection using Hybrid Network Models. arXiv:2510.10321
- Gajjar et al. (2025b) Jugal Gajjar, Kamalasankari Subramaniakuppusamy, and Noha El Kachach. 2025b. MalCodeAI: Autonomous Vulnerability Detection and Remediation via Language Agnostic Code Reasoning. In IEEE International Conference on Information Reuse and Integration and Data Science (IRI). IEEE, San Jose, CA, USA, 31–36. doi:10.1109/IRI66576.2025.00014
- Gao et al. (2023) Shuzheng Gao, Hongyu Zhang, Cuiyun Gao, and Chaozheng Wang. 2023. Keeping Pace with Ever-Increasing Data: Towards Continual Learning of Code Intelligence Models. In International Conference on Software Engineering (ICSE). IEEE/ACM, Melbourne, Australia, 30–42. doi:10.1109/ICSE48619.2023.00015
@@
- Germano et al. (2025) Lucas Bastos Germano, Ronaldo Ribeiro Goldschmidt, Ricardo Choren Noya, and Julio Cesar Duarte. 2025. A Systematic Review on Detection, Repair, and Explanation of Vulnerabilities in Source Code Using Large Language Models. IEEE Access 13 (2025), 192263–192293. doi:10.1109/ACCESS.2025.3631363
- Gnieciak and Szandala (2025) Damian Gnieciak and Tomasz Szandala. 2025. Large Language Models Versus Static Code Analysis Tools: A Systematic Benchmark for Vulnerability Detection. arXiv:2508.04448